# Orthogonal Howe duality and dynamical (split) symmetric pairs

### v2

@@ 10.2. Representation
Recall the spinor module $\mathfrak{P}_{nm}$ from Section 10.1.3. The action of $\mathrm{Cl}_{nm}$ on $\mathfrak{P}_{nm}$ gives rise to a total action of $\mathfrak{so}_{2n}$ of $\mathfrak{P}_{nm}$ by restricting along the coproduct:
@@
On the $\mathrm{O}_{m}$-side, we first consider $\mathfrak{P}_{nm}=\Lambda^{\bullet}(\mathbb{C}^{n}\otimes\mathbb{C}^{m})$ as a $\mathrm{GL}_{m}$-representation, see Section 9.2. In particular, we have the embeddings
@@
and the total $\mathfrak{gl}_{m}$ action is given by the sum $E_{\alpha\beta}\mapsto\sum_{i}E_{\alpha\beta}^{(i)}$. As in Section 10.1.2, we denote by $\theta$ the Cartan involution, $\mathrm{O}_{m}:=\mathrm{GL}_{m}^{\theta}$ the fixed point subgroup, and by $B_{\alpha\beta}=E_{\beta\alpha}-E_{\alpha\beta}$ the basis of the fixed point Lie subalgebra $\mathfrak{so}_{m}:=\mathfrak{gl}_{m}^{\theta}$.
One can easily see that the actions of $\mathfrak{so}_{2n}$ and $\mathrm{O}_{m}$ commute. In fact, there is an analog of Theorem 9.1, for instance see [ChengWang, Corollary 5.41]. The irreducible highest-weight $\mathfrak{so}_{2n}$-modules below are with respect to the following choice of the Borel subalgebra:
$\epsilon_{1}-\epsilon_{2}$$-\epsilon_{1}-\epsilon_{2}$$\epsilon_{2}-\epsilon_{3}$$\epsilon_{3}-\epsilon_{4}$$\epsilon_{n-1}-\epsilon_{n}$
@@
Observe that we have an isomorphism $\mathrm{Cl}_{nm}\cong\mathrm{Cl}^{(1)}_{m}\otimes\ldots\otimes\mathrm{Cl}^{(n)}_{m}$ of superalgebras. Denote by $\tau_{j}$ the Fourier transform (10.4) along the $j$-th component. One can easily see that under the embedding $\mathfrak{so}_{2n}\hookrightarrow\mathrm{Cl}_{nm}$, it corresponds to the Dynkin automorphism of $\mathfrak{so}_{2n}$ from (10.1). On the $\mathrm{O}_{m}$-side, it corresponds to a shifted version $\tau$ of the Cartan involution from Section 2.4:
@@ 10.3. Classical boundary duality
One can see that these operators define a connection gauge equivalent (in the sense of 8.3) to the Casimir connection of [FelderMarkovTarasovVarchenko] for $\mathfrak{so}_{2n}$.
@@
Here, $\Omega_{\mathfrak{gl}_{n}}^{(i,j)}$ is the split Casimir operator as in Section 9 and $\Omega^{(i)}_{\mathfrak{so}_{m}}$ is the Casimir element $\Omega_{\mathfrak{so}_{m}}=\frac{1}{2}\sum_{\alpha<\beta}B_{\alpha\beta}^{2}$.
@@ 10.4. Classical duality
$\Omega_{\mathfrak{so}_{2n}}^{(\alpha,\beta)}=\frac{1}{2}\sum_{i,j=1}^{n}(M_{ij}^{(\alpha)}M_{ji}^{(\beta)}+M_{i,-j}^{(\alpha)}M_{-j,i}^{(\beta)}+M_{-j,i}^{(\alpha)}M_{i,-j}^{(\beta)}).$
@@
(10.8) $\nabla^{\mathrm{KZ},\langle n\rangle}_{z_{\alpha}}=\kappa\cdot\partial_{z_{\alpha}}-\sum_{\beta\neq\alpha}\frac{\Omega_{\mathfrak{so}_{2n}}^{(\alpha,\beta)}-n}{z_{\alpha}-z_{\beta}}.$
On the $\mathrm{O}_{m}$-side, consider the differential boundary dynamical operators of 8.6
@@
(10.9) $\nabla^{\mathrm{bCas},\langle m\rangle}_{\lambda_{\alpha}}(z,\kappa)=\kappa\cdot\partial_{\lambda_{\alpha}}-\sum_{\beta\neq\alpha}\frac{B_{\alpha\beta}^{2}}{z_{\alpha}-z_{\beta}}.$
Identifying $\mathbb{C}^{m}\rightarrow\mathfrak{h}_{\mathfrak{gl}_{m}}$ by $z_{\alpha}\mapsto h_{\alpha}$, the following proposition can be proved analogously to Theorem 10.8.
@@ 10.5. Quantum duality
In this subsection, we show the difference analog of Theorem 10.8 and Theorem 10.9.
@@ 10.5.1. Yangian
As in Section 9.4.1, there is an intertwining operator ˇ$\check{\mathcal{R}}_{12}(z_{1}-z_{2})\colon\mathfrak{P}_{n}(z_{1})\otimes\mathfrak{P}_{n}(z_{2})\rightarrow\mathfrak{P}(z_{2})\otimes\mathfrak{P}(z_{1})$ that we call the $R$-operator (as opposed to an $R$-matrix which does not permute factors). It satisfies
(10.10) ˇ$\check{\mathcal{R}}_{12}(u)(M_{ij}^{\mathfrak{P},(1)}+M_{ij}^{\mathfrak{P},(2)})$ $=(M_{ij}^{\mathfrak{P},(1)}+M_{ij}^{\mathfrak{P},(2)})\check{\mathcal{R}}_{12}(u),$
(10.11) ˇ$\check{\mathcal{R}}_{12}(u)(-uM_{ij}^{\mathfrak{P},(2)}+\sum_{k=-n}^{n}M_{ik}^{\mathfrak{P},(1)}M_{kj}^{\mathfrak{P},(2)})$ $=(-uM_{ij}^{\mathfrak{P},(1)}+\sum_{k=-n}^{n}M_{ik}^{\mathfrak{P},(1)}M_{kj}^{\mathfrak{P},(2)})\check{\mathcal{R}}_{12}(u).$
By the $(\mathfrak{so}_{2n},\mathrm{O}_{2})$-duality of Theorem 10.6, it follows from (10.10) that ˇ$\check{\mathcal{R}}_{12}(u)$ depends only on
@@
We indicate this dependence by ˇ$\check{\mathcal{R}}_{12}(u)=\check{\mathcal{R}}_{12}(u|b^{\langle 2\rangle}_{12})$. Observe that by the duality, the operator $b^{\langle 2\rangle}_{12}$ is diagonalizable with integral eigenvalues.
@@
ˇ$\check{\mathcal{R}}_{12}(u|b^{\langle 2\rangle}_{12})=r(u)\cdot\mathrm{B}\left(\frac{ib^{\langle 2\rangle}_{12}+1+u}{2},-u\right)^{-1}$
@@
ˇ$\check{\mathcal{R}}_{12}(t)=\frac{(-1)^{\deg^{(1)}\deg^{(2)}+1}(-2i)^{-t}}{t}\cdot\mathrm{B}\left(\frac{ib^{\langle 2\rangle}_{12}+1+t}{2},-t\right)^{-1}.$
@@ 10.5.2. Twisted Yangian
Consider the universal reflection equation (10.13) acting on the tensor product $V_{1}(z_{1})\otimes V_{2}(z_{2})$ of $\mathrm{Y}(\mathfrak{gl}_{m})$-representations. Since the universal $S$-matrix commutes with $\mathfrak{so}_{m}$, we conclude by Proposition 10.12 and irreducibility of $\mathfrak{so}_{m}$-modules $V_{1}$ and $V_{2}$ that the universal $S$-matrices act by scalar functions. Therefore, the reflection equation is reduced to
@@
Observe that the action of $(\tilde{\theta}\otimes\tilde{\theta})\mathfrak{R}$ on $V_{1}(z_{1})\otimes V_{2}(z_{2})$ satisfies the properties (9.10) for the permuted $R$-matrix $\mathcal{R}_{V_{2}V_{1}}(t)$ except for, possibly, the normalization condition. In particular, it is proportional to $\mathcal{R}_{V_{2}V_{1}}(z_{2}-z_{1})$. By unitarity (9.11), the latter is equal to $\mathcal{R}_{V_{1}V_{2}}(z_{1}-z_{2})^{-1}$. Likewise, the operator $(1\otimes\tilde{\theta})\mathfrak{R}$ acting on $V_{1}(z_{1})\otimes V_{2}(z_{2})$ satisfies the properties (10.14) up to normalization. Therefore, applying $\tilde{\theta}\otimes\tilde{\theta}$ to both sides of (10.15) and rearranging the terms, we have
@@ 10.5.3. Duality
Recall from Section 10.1.1 that the positive roots of $\mathfrak{so}_{2n}$ are of the form $\epsilon_{i}\pm\epsilon_{j}$ for $0<i<j$. Consider the root $\epsilon_{i}-\epsilon_{j}$. We have
@@
By construction, they come from the embedding $\mathfrak{sl}_{n}\hookrightarrow\mathfrak{gl}_{n}\hookrightarrow\mathfrak{so}_{2n}$. Recall from Section 10.1.2 that we have an isomorphism $\mathfrak{P}_{nm}\cong\Lambda^{\bullet}(\mathbb{C}^{n}\otimes\mathbb{C}^{m})$ of $\mathfrak{sl}_{n}$-representations.
Recall the dynamical $B$-operators (6.3). Because of the isomorphism, we see that the action of the $B$-operators $\mathcal{B}_{\epsilon_{i}-\epsilon_{j}}(t)$ for $\mathfrak{so}_{2n}$ coincides with those for $\mathfrak{gl}_{n}$ (since they depend only on the $\mathfrak{sl}_{n}$-structure). On the $\mathrm{O}_{m}$-side, consider the $R$-matrix $\mathcal{R}_{ij}^{\langle m\rangle}(t)$ for $\mathfrak{gl}_{m}$ as in Section 9.4.1. Denote by
@@
The following proposition is a tautological reformulation of Theorem 9.1.
@@
Consider a positive root $\epsilon_{i}+\epsilon_{j}$ for $i<j$. Recall the involution $\tau_{j}$ of (10.5). As it was discussed in Section 10.2, it corresponds to (a version of) the Dynkin automorphism of $\mathfrak{so}_{2n}$, so,
@@
in particular, $\tau_{j}(\mathcal{B}_{\epsilon_{i}-\epsilon_{j}}^{\langle n\rangle}(t))=\mathcal{B}_{\epsilon_{i}+\epsilon_{j}}^{\langle n\rangle}(t)$. Also, recall the automorphism $\tau_{j}^{\mathfrak{P}}$ of $\mathfrak{P}_{nm}$ from (10.5). It follows from Proposition 10.5 that $\mathrm{Ad}_{\tau_{j}^{\mathfrak{P}}}(\Gamma)=\tau_{j}(\Gamma)$ for any $\Gamma\in\mathrm{Cl}_{nm}$. In particular, we have $\mathrm{Ad}_{\tau_{j}^{\mathfrak{P}}}(\mathcal{B}_{\epsilon_{i}-\epsilon_{j}}^{\langle n\rangle}(t))=\mathcal{B}_{\epsilon_{i}+\epsilon_{j}}^{\langle n\rangle}(t)$.
@@
the $\tau_{j}$-twisted factor (10.17). The following proposition follows directly from Proposition 10.14 by applying $\mathrm{Ad}_{\tau^{\mathfrak{P}}_{j}}$ on both sides.
@@
$[R^{--}_{ij}(t),R_{ij}^{-+}(s)]=0\ \mathrm{for\ any}\ s,t$
is satisfied as well: for $i<j$, observe that any exterior power $\Lambda^{k}\mathbb{C}^{m}$ is an irreducible representation of $\mathfrak{so}_{m}$, therefore, the version of the reflection equation of Proposition 10.13 gives the commutation relation for a suitable choices of parameters $z_{1},z_{2}$. Likewise, for $i>j$, we have $R_{ij}^{-+}(t)=\mathcal{R}^{t_{2},\langle m\rangle}_{ji}(t)$ and $R^{--}_{ij}(t)=\mathcal{R}^{\langle m\rangle}_{ij}(t)=\mathcal{R}^{\langle m\rangle}_{ji}(-t)^{-1}$ by the unitarity relation. So, we need to show that
@@
For a suitable choices of parameters $z_{1},z_{2}$, this is equivalent to the reflection equation of Proposition 10.13.
@@
For a collection of variables $z=(z_{1},\ldots,z_{n})$ and $\alpha\in\mathbb{C}$, denote by $z+\alpha:=(z_{1}+\alpha,\ldots,z_{n}+\alpha)$. Then, Proposition 10.16 implies the following result.
@@
Finally, consider the spinorial $R$-operator ˇ$\check{\mathcal{R}}^{\langle n\rangle}_{\alpha\beta}(t)$ of 10.11 acting on the $(\alpha,\beta)$-components of $\mathfrak{P}_{nm}$, and by $\mathcal{R}_{\alpha\beta}^{\langle n\rangle}(t)=P_{\alpha\beta}\check{\mathcal{R}}^{\langle n\rangle}_{\alpha\beta}(t)$ the corresponding $R$-matrix. Recall the boundary dynamical Weyl group operator $\mathcal{A}^{\langle m\rangle}_{\epsilon_{\alpha}-\epsilon_{\beta}}(t):=A^{\mathfrak{so}_{m},\epsilon_{\alpha}-\epsilon_{\beta}}_{\mathfrak{P}_{nm}}(t)$ of 7.15 and the corresponding $B$-operator of 7.19, which we denote by $\mathcal{B}^{\langle m\rangle}_{\epsilon_{\alpha}-\epsilon_{\beta}}(t)$.
@@
where $\tilde{s}_{12}^{\mathfrak{so}_{2}}$ is the boundary Weyl group operator of 7.19. Let $\deg_{\alpha}$ be the degree operator acting on the $\alpha$-component; one can check that $\tilde{s}^{\mathfrak{so}_{2}}_{12}=P_{12}(-1)^{\deg_{1}\deg_{2}+\deg_{2}}$. Indeed: consider a monomial $x_{i_{1},1}\ldots x_{i_{k},1}x_{j_{1},2}\ldots x_{j_{l},2}$. By (3.2) and tensor property of $\tilde{s}_{12}^{\mathfrak{so}_{2}}$, we have
@@
By 10.11 and the construction of the $A$-operator, we have
ˇ$\check{\mathcal{R}}_{12}^{\langle n\rangle}(t)=(-1)^{\deg_{1}\deg_{2}+1}\mathcal{A}_{\epsilon_{1}-\epsilon_{2}}^{\langle 2\rangle}(-t-1).$
@@
Analogously to Section 9.4.2, define
@@
By [CherednikQKZ], both collections of operators are consistent. The following theorem is a boundary analog of Theorem 9.4.

### v3

@@ 10.2. Representation
Recall the spinor module $\mathfrak{P}_{nm}$ from Section˜10.1.3. The action of $\mathrm{Cl}_{nm}$ on $\mathfrak{P}_{nm}$ gives rise to a total action of $\mathfrak{so}_{2n}$ of $\mathfrak{P}_{nm}$ by restricting along the coproduct:
@@
On the $\mathrm{O}_{m}$-side, we first consider $\mathfrak{P}_{nm}=\Lambda^{\bullet}(\mathbb{C}^{n}\otimes\mathbb{C}^{m})$ as a $\mathrm{GL}_{m}$-representation, see Section˜9.2. In particular, we have the embeddings
@@
and the total $\mathfrak{gl}_{m}$ action is given by the sum $E_{\alpha\beta}\mapsto\sum_{i}E_{\alpha\beta}^{(i)}$. As in Section˜10.1.2, we denote by $\theta$ the Cartan involution, $\mathrm{O}_{m}:=\mathrm{GL}_{m}^{\theta}$ the fixed point subgroup, and by $B_{\alpha\beta}=E_{\beta\alpha}-E_{\alpha\beta}$ the basis of the fixed point Lie subalgebra $\mathfrak{so}_{m}:=\mathfrak{gl}_{m}^{\theta}$.
One can easily see that the actions of $\mathfrak{so}_{2n}$ and $\mathrm{O}_{m}$ commute. In fact, there is an analog of Theorem˜9.1, for instance see [ChengWang, Corollary 5.41]. The irreducible highest-weight $\mathfrak{so}_{2n}$-modules below are with respect to the following choice of the Borel subalgebra:
$\epsilon_{1}-\epsilon_{2}$$\epsilon_{1}-\epsilon_{2}$$\epsilon_{2}-\epsilon_{3}$$\epsilon_{3}-\epsilon_{4}$$\epsilon_{n-1}-\epsilon_{n}$
@@
Observe that we have an isomorphism $\mathrm{Cl}_{nm}\cong\mathrm{Cl}^{(1)}_{m}\otimes\ldots\otimes\mathrm{Cl}^{(n)}_{m}$ of superalgebras. Denote by $\tau_{j}$ the Fourier transform (10.4) along the $j$-th component. One can easily see that under the embedding $\mathfrak{so}_{2n}\hookrightarrow\mathrm{Cl}_{nm}$, it corresponds to the Dynkin automorphism of $\mathfrak{so}_{2n}$ from (10.1). On the $\mathrm{O}_{m}$-side, it corresponds to a shifted version $\tau$ of the Cartan involution from Section˜2.4:
@@ 10.3. Classical boundary duality
One can see that these operators define a connection gauge equivalent (in the sense of ˜8.3) to the Casimir connection of [FelderMarkovTarasovVarchenko] for $\mathfrak{so}_{2n}$.
@@
Here, $\Omega_{\mathfrak{gl}_{n}}^{(i,j)}$ is the split Casimir operator as in Section˜9 and $\Omega^{(i)}_{\mathfrak{so}_{m}}$ is the Casimir element $\Omega_{\mathfrak{so}_{m}}=\frac{1}{2}\sum_{\alpha<\beta}B_{\alpha\beta}^{2}$.
@@ 10.4. Classical duality
$\Omega_{\mathfrak{so}_{2n}}^{(\alpha,\beta)}=\frac{1}{2}\sum_{i,j=1}^{n}M_{ij}^{(\alpha)}M_{ji}^{(\beta)}+\frac{1}{2}\sum_{i<j}(M_{i,-j}^{(\alpha)}M_{-j,i}^{(\beta)}+M_{-j,i}^{(\alpha)}M_{i,-j}^{(\beta)}).$
@@
(10.8) $\nabla^{\mathrm{KZ},\langle n\rangle}_{z_{\alpha}}=\kappa\cdot\partial_{z_{\alpha}}-\sum_{\beta\neq\alpha}\frac{\Omega_{\mathfrak{so}_{2n}}^{(\alpha,\beta)}-n/8}{z_{\alpha}-z_{\beta}}.$
On the $\mathrm{O}_{m}$-side, consider a slight modification of the differential boundary dynamical operators of ˜8.6
@@
(10.9) $\nabla^{\mathrm{bCas},\langle m\rangle}_{\lambda_{\alpha}}(z,\kappa)=\kappa\cdot\partial_{\lambda_{\alpha}}+\frac{1}{4}\sum_{\beta\neq\alpha}\frac{B_{\alpha\beta}^{2}}{z_{\alpha}-z_{\beta}}.$
Identifying $\mathbb{C}^{m}\rightarrow\mathfrak{h}_{\mathfrak{gl}_{m}}$ by $z_{\alpha}\mapsto h_{\alpha}$, the following proposition can be proved analogously to Theorem˜10.8.
@@ 10.5. Quantum duality
In this subsection, we show the difference analog of Theorem˜10.8 and Theorem˜10.9.
@@ 10.5.1. Yangian
As in Section˜9.4.1, there is an intertwining operator $\check{\mathcal{R}}_{12}(z_{1}-z_{2})\colon\mathfrak{P}_{n}(z_{1})\otimes\mathfrak{P}_{n}(z_{2})\rightarrow\mathfrak{P}(z_{2})\otimes\mathfrak{P}(z_{1})$ that we call the $R$-operator (as opposed to an $R$-matrix which does not permute factors). It satisfies
(10.10) $\check{\mathcal{R}}_{12}(u)(M_{ij}^{\mathfrak{P},(1)}+M_{ij}^{\mathfrak{P},(2)})$ $=(M_{ij}^{\mathfrak{P},(1)}+M_{ij}^{\mathfrak{P},(2)})\check{\mathcal{R}}_{12}(u),$
(10.11) $\check{\mathcal{R}}_{12}(u)(-uM_{ij}^{\mathfrak{P},(2)}+\sum_{k=-n}^{n}M_{ik}^{\mathfrak{P},(1)}M_{kj}^{\mathfrak{P},(2)})$ $=(-uM_{ij}^{\mathfrak{P},(1)}+\sum_{k=-n}^{n}M_{ik}^{\mathfrak{P},(1)}M_{kj}^{\mathfrak{P},(2)})\check{\mathcal{R}}_{12}(u).$
By the $(\mathfrak{so}_{2n},\mathrm{O}_{2})$-duality of Theorem˜10.6, it follows from (10.10) that $\check{\mathcal{R}}_{12}(u)$ depends only on
@@
We indicate this dependence by $\check{\mathcal{R}}_{12}(u)=\check{\mathcal{R}}_{12}(u|b^{\langle 2\rangle}_{12})$. Observe that by the duality, the operator $b^{\langle 2\rangle}_{12}$ is diagonalizable with integral eigenvalues.
@@
$\check{\mathcal{R}}_{12}(u|b^{\langle 2\rangle}_{12})=r(u)\cdot\mathrm{B}\left(\frac{ib^{\langle 2\rangle}_{12}+1+u}{2},-u\right)^{-1}$
@@
$\check{\mathcal{R}}_{12}(t)=\frac{(-1)^{\deg^{(1)}\deg^{(2)}+1}(-2i)^{-t}}{t}\cdot\mathrm{B}\left(\frac{ib^{\langle 2\rangle}_{12}+1+t}{2},-t\right)^{-1}.$
@@ 10.5.2. Twisted Yangian
Consider the universal reflection equation (10.13) acting on the tensor product $V_{1}(z_{1})\otimes V_{2}(z_{2})$ of $\mathrm{Y}(\mathfrak{gl}_{m})$-representations. Since the universal $S$-matrix commutes with $\mathfrak{so}_{m}$, we conclude by Proposition˜10.12 and irreducibility of $\mathfrak{so}_{m}$-modules $V_{1}$ and $V_{2}$ that the universal $S$-matrices act by scalar functions. Therefore, the reflection equation is reduced to
@@
[…]
@@ 10.5.3. Duality
Recall from Section˜10.1.1 that the positive roots of $\mathfrak{so}_{2n}$ are of the form $\epsilon_{i}\pm\epsilon_{j}$ for $0<i<j$. Consider the root $\epsilon_{i}-\epsilon_{j}$. We have
@@
By construction, they come from the embedding $\mathfrak{sl}_{n}\hookrightarrow\mathfrak{gl}_{n}\hookrightarrow\mathfrak{so}_{2n}$. Recall from Section˜10.1.2 that we have an isomorphism $\mathfrak{P}_{nm}\cong\Lambda^{\bullet}(\mathbb{C}^{n}\otimes\mathbb{C}^{m})$ of $\mathfrak{sl}_{n}$-representations.
Recall the dynamical $B$-operators (6.3). Because of the isomorphism, we see that the action of the $B$-operators $\mathcal{B}_{\epsilon_{i}-\epsilon_{j}}(t)$ for $\mathfrak{so}_{2n}$ coincides with those for $\mathfrak{gl}_{n}$ (since they depend only on the $\mathfrak{sl}_{n}$-structure). On the $\mathrm{O}_{m}$-side, consider the $R$-matrix $\mathcal{R}_{ij}^{\langle m\rangle}(t)$ for $\mathfrak{gl}_{m}$ as in Section˜9.4.1. Denote by
@@
The following proposition is a tautological reformulation of Theorem˜9.1.
@@
Consider a positive root $\epsilon_{i}+\epsilon_{j}$ for $i<j$. Recall the involution $\tau_{j}$ of (10.5). As it was discussed in Section˜10.2, it corresponds to (a version of) the Dynkin automorphism of $\mathfrak{so}_{2n}$, so,
@@
in particular, $\tau_{j}(\mathcal{B}_{\epsilon_{i}-\epsilon_{j}}^{\langle n\rangle}(t))=\mathcal{B}_{\epsilon_{i}+\epsilon_{j}}^{\langle n\rangle}(t)$. Also, recall the automorphism $\tau_{j}^{\mathfrak{P}}$ of $\mathfrak{P}_{nm}$ from (10.5). It follows from Proposition˜10.5 that $\mathrm{Ad}_{\tau_{j}^{\mathfrak{P}}}(\Gamma)=\tau_{j}(\Gamma)$ for any $\Gamma\in\mathrm{Cl}_{nm}$. In particular, we have $\mathrm{Ad}_{\tau_{j}^{\mathfrak{P}}}(\mathcal{B}_{\epsilon_{i}-\epsilon_{j}}^{\langle n\rangle}(t))=\mathcal{B}_{\epsilon_{i}+\epsilon_{j}}^{\langle n\rangle}(t)$.
@@
the $\tau_{j}$-twisted factor (10.17). The following proposition follows directly from Proposition˜10.14 by applying $\mathrm{Ad}_{\tau^{\mathfrak{P}}_{j}}$ on both sides.
@@
$R^{--}_{ij}(t),R_{ij}^{-+}(s)]=0\ \mathrm{for\ any}\ s,t$
is satisfied as well: for $i<j$, observe that any exterior power $\Lambda^{k}\mathbb{C}^{m}$ is an irreducible representation of $\mathfrak{so}_{m}$, therefore, the version of the reflection equation of Proposition˜10.13 gives the commutation relation for a suitable choices of parameters $z_{1},z_{2}$. Likewise, for $i>j$, we have $R_{ij}^{-+}(t)=\mathcal{R}^{t_{2},\langle m\rangle}_{ji}(t)$ and $R^{--}_{ij}(t)=\mathcal{R}^{\langle m\rangle}_{ij}(t)=\mathcal{R}^{\langle m\rangle}_{ji}(-t)^{-1}$ by the unitarity relation. So, we need to show that
@@
For a suitable choices of parameters $z_{1},z_{2}$, this is equivalent to the reflection equation of Proposition˜10.13.
@@
For a collection of variables $z=(z_{1},\ldots,z_{n})$ and $\alpha\in\mathbb{C}$, denote by $z+\alpha:=(z_{1}+\alpha,\ldots,z_{n}+\alpha)$. Then, Proposition˜10.16 implies the following result.
@@
Finally, consider the spinorial $R$-operator $\check{\mathcal{R}}^{\langle n\rangle}_{\alpha\beta}(t)$ of ˜10.11 acting on the $(\alpha,\beta)$-components of $\mathfrak{P}_{nm}$, and by $\mathcal{R}_{\alpha\beta}^{\langle n\rangle}(t)=P_{\alpha\beta}\check{\mathcal{R}}^{\langle n\rangle}_{\alpha\beta}(t)$ the corresponding $R$-matrix. Recall the boundary dynamical Weyl group operator $\mathcal{A}^{\langle m\rangle}_{\epsilon_{\alpha}-\epsilon_{\beta}}(t):=A^{\mathfrak{so}_{m},\epsilon_{\alpha}-\epsilon_{\beta}}_{\mathfrak{P}_{nm}}(t)$ of ˜7.15 and the corresponding $B$-operator of ˜7.19, which we denote by $\mathcal{B}^{\langle m\rangle}_{\epsilon_{\alpha}-\epsilon_{\beta}}(t)$.
@@
where $\tilde{s}_{12}^{\mathfrak{so}_{2}}$ is the boundary Weyl group operator of ˜7.19. Let $\deg_{\alpha}$ be the degree operator acting on the $\alpha$-component; one can check that $\tilde{s}^{\mathfrak{so}_{2}}_{12}=P_{12}(-1)^{\deg_{1}\deg_{2}+\deg_{2}}$. Indeed: consider a monomial $x_{i_{1},1}\ldots x_{i_{k},1}x_{j_{1},2}\ldots x_{j_{l},2}$. By (3.2) and tensor property of $\tilde{s}_{12}^{\mathfrak{so}_{2}}$, we have
@@
By ˜10.11 and the construction of the $A$-operator, we have
$\check{\mathcal{R}}_{12}^{\langle n\rangle}(t)=(-1)^{\deg_{1}\deg_{2}+1}\mathcal{A}_{\epsilon_{1}-\epsilon_{2}}^{\langle 2\rangle}(-t-1).$
@@
Analogously to Section˜9.4.2, define
@@
By [CherednikQKZ], both collections of operators are consistent. The following theorem is a boundary analog of Theorem˜9.4.